Q64 / 150 · MHT CET 2026 (15 April Shift 2) · Mathematics
The probability that event A happens in a trial is 0.4. If three independent trials are made, then the probability that A happens at least once is.......
- A$0.216$
- B$0.926$
- C$0.064$
- Correct answerD$0.784$
Solution
Let $P(A)$ be the probability that event A happens in a single trial.
$P(A) = 0.4$
The probability that event A does not happen in a single trial is $P(A') = 1 - 0.4 = 0.6$.
Since the trials are independent, the probability that A does not happen in any of the three trials is $(P(A'))^3 = (0.6)^3 = 0.216$.
The probability that A happens at least once in three trials is $1 - P(\text{A does not happen in any trial})$.
$1 - 0.216 = 0.784$
Answer: $0.784$
$P(A) = 0.4$
The probability that event A does not happen in a single trial is $P(A') = 1 - 0.4 = 0.6$.
Since the trials are independent, the probability that A does not happen in any of the three trials is $(P(A'))^3 = (0.6)^3 = 0.216$.
The probability that A happens at least once in three trials is $1 - P(\text{A does not happen in any trial})$.
$1 - 0.216 = 0.784$
Answer: $0.784$
Hence, the correct answer is option: 'D'.